: Payments of $ 670 are being made at the end of each month for 5 years at an interest of 8% compounded monthly. Calculate the Present Value.
1
Expert's answer
2019-11-20T13:13:53-0500

Calculate the Present Value


We need to find the present value


Solution:


Given,


P = Principal value = $670


t = 5 years


r = 8% = 0.08


n = number of time interest compounded = 12


A = Amount


W e know the Formula for present value ,


Present Value of Annuity

"= P \times [1 - ( 1 + \frac {r}{n}) ^ {- t \times n }] \times [ \frac {1 + \frac {r}{n}} {(\frac {r} {n})}]"


"= 670 \times [1 - ( 1 + \frac {0.08}{12}) ^ {- 5 \times 12 }] \times [ \frac {1 + \frac {0.08}{12}} {(\frac {0.08} {12})}]"

"= 670 \times [1 - ( 1 + {0.0066}) ^ {- 60}] \times [ \frac {1 +0.0066} {0.0066}]"


"= 670 \times 0.326 \times 152.51 = \$ 33,311.23"




Answer : Present value = $33,311.23